# Two-branch retention behavior in unsaturated fractured rock driven by fracture–matrix flow partitioning

### 27 Mar 2026 v1

@@ 1 Introduction
Unsaturated flow in fractured rock plays a critical role in subsurface fluid transport, with important implications for groundwater management and nuclear waste disposal (Pruess & Tsang, 1990; Liu et al., 1998; Nimmo, 2012; Rutqvist & Tsang, 2012; Zhang, 2018; Wang et al., 2019; Liang et al., 2022). However, predicting unsaturated flow in such systems remains challenging, because fractured media consist of two distinct domains, i.e., the rock matrix and the fracture network, with strongly contrasting hydraulic and retention properties (Glass et al., 1995; Seol et al., 2003; Huang et al., 2005). Unlike saturated conditions, where fractures typically dominate fluid flow, under unsaturated conditions, the relative importance of matrix versus fracture flow remains unresolved.
From a theoretical perspective, the porous rock matrix is expected to play a dominant role under unsaturated conditions, because strong capillary forces promote imbibition into the fine pore space of the matrix, which can suppress rapid flow through fractures and lead to relatively diffuse infiltration (Wang & Narasimhan, 1985; Nitao & Buscheck, 1991). Based on this conceptual view, subsurface flow has been traditionally represented using continuum approaches that treat fractured rock as an equivalent porous medium (Peters & Klavetter, 1988). However, field observations have revealed a different behavior. Field experiments at sites such as Yucca Mountain (Nevada, United States) have shown that water can migrate rapidly through localized preferential pathways along fractures even under unsaturated conditions (Pruess, 1998; Pruess et al., 1999; Bodvarsson et al., 2003). Laboratory experiments have also illustrated that flow can proceed with minimal imbibition into the surrounding rock matrix, thereby limiting fracture–matrix hydraulic exchange and enabling fast fracture-dominated transport (Liu et al., 1998; Wood et al., 2002; Ji et al., 2004; Tokunaga et al., 2005; Zhou et al., 2022). Furthermore, the geometry and aperture distribution of fracture networks strongly influence recharge efficiency and the rate at which water percolates through the unsaturated zone (Liu et al., 2002; Cey et al., 2006; Zhu et al., 2021).
@@ 3.1 Emergent two-branch retention behavior
$p_{c}(S_{e})=\begin{cases}p_{cc}\left(\dfrac{S_{e}}{S_{c}}\right)^{-\frac{1}{n_{m}}},&S_{e}\leq S_{c}\\[7.5pt] p_{cc}-\left(\dfrac{\rho g}{\alpha_{f}}+p_{cc}\right)\left(\dfrac{S_{e}-S_{c}}{1-S_{c}}\right)^{-\frac{1}{n_{f}}},&S_{e}>S_{c}\end{cases}$ (12)
where $k_{rc}$ is the relative permeability at $S_{c}$ and $p_{cc}$ is the capillary pressure at $S_{c}$. This generalized formulation extends the classical Brooks-Corey relationship by introducing regime-dependent normalization and incorporating the critical saturation (associated with $k_{rc}$ and $p_{cc}$) separating the two regimes. If the effective saturation $S_{e}$ is below the critical saturation $S_{c}$, the system resides in the matrix-dominated regime, in which the constitutive relationships are expressed in terms of the normalized saturation $S_{e}/S_{c}$. For $S_{e}>S_{c}$, corresponding to the fracture-dominated regime, where the constitutive relationships are expressed in terms of the normalized critical saturation $(S_{e}-S_{c})/(1-S_{c})$. The constitutive parameters $n$ and $l$ are assigned with matrix properties for $S_{e}$ below $S_{c}$, and assigned with fracture properties for $S_{e}$ above $S_{c}$. In the limits $S_{c}\to 0$ and $S_{c}\to 1$, the generalized formulation reduces to the classical Brooks–Corey model for fracture-only and matrix-only flow, respectively. This two-branch analytical formulation closely matches the numerical simulation results and well captures the regime transition at $S_{c}$ (Figure 3).
@@ 3.2 Transition between matrix- and fracture-dominated regimes
$h_{c}=\left[\frac{k_{f}\int A_{f}(L)\,dL}{k_{m}\int A_{m}(L)\,dL}\frac{\alpha_{m}^{2+n_{m}(l_{m}+2)}}{\alpha_{f}^{2+n_{f}(l_{f}+2)}}\right]^{\frac{1}{n_{f}(l_{f}+2)-n_{m}(l_{m}+2)}}$ (15)
@@
Here, we use the approximation $(1-b\gamma)\approx 1$ given that $b\gamma\ll 1$. Combining Equations 9 and 16, and invoking the cubic law for fracture permeability, the critical pressure head $h_{c}$ can be expressed explicitly as a function of fracture aperture:
@@ 4.2 Control of fracture network connectivity
The two-branch retention behavior persists across fracture networks with varying connectivity. Re-examination of Equation 16 shows that the critical transition is controlled primarily by the fracture intensity. The percolation parameter quantifying fracture network connectivity does not explicitly appear in the formulation, but it influences the validity of the analytical prediction. As shown in Figure 4, the analytical solution systematically underestimates the dimensionless critical pressure head for disconnected networks. This discrepancy arises from the assumption in Equation 16 that fractures form fully connected pathways for flow. While this assumption may hold for well-connected networks, where multiple flow paths distribute water flux, it breaks down near or below the percolation threshold (Davy et al., 2024). In disconnected networks, fractures alone cannot transmit flow through the system, and a substantial fraction of the flux must pass through matrix, thereby involving fracture–matrix exchange (Yoon et al., 2023), consistent with the limited increase in relative permeability between the critical saturation and full saturation (Text S2 and Figure S8 in Supporting Information). The fracture intensity used in Equation 16 therefore overestimates the hydraulically active fracture population in disconnected networks, leading to an underestimation of the dimensionless critical pressure head.
@@ Acknowledgments
Derivations, benchmark simulations, sensitivity analyses for both single-fracture and fracture network cases, comparisons between numerical and analytical results, and additional evaluations of hydraulic parameter effects and retention models are provided in the Supporting Information.

### 01 Apr 2026 v2

@@ 1 Introduction
Unsaturated flow in fractured rock plays a critical role in subsurface fluid transport, with important implications for groundwater management and nuclear waste disposal (Pruess & Tsang, 1990; Liu et al., 1998; Nimmo, 2012; Rutqvist & Tsang, 2012; Zhang, 2018; Wang et al., 2019; Liang et al., 2022; Viswanathan et al., 2022). However, predicting unsaturated flow in such systems remains challenging, because fractured media consist of two distinct domains, i.e., the rock matrix and the fracture network, with strongly contrasting hydraulic and retention properties (Glass et al., 1995; Seol et al., 2003; Huang et al., 2005). Unlike saturated conditions, where fractures typically dominate fluid flow, under unsaturated conditions, the relative importance of matrix versus fracture flow remains unresolved.
From a theoretical perspective, the porous rock matrix is expected to play a dominant role under unsaturated conditions, because strong capillary forces promote imbibition into the fine pore space of the matrix, which can suppress rapid flow through fractures and lead to relatively diffuse infiltration (Wang & Narasimhan, 1985; Nitao & Buscheck, 1991). Based on this conceptual view, subsurface flow has been traditionally represented using continuum approaches that treat fractured rock as an equivalent porous medium (Peters & Klavetter, 1988). However, field observations have revealed a different behavior. Field experiments at sites such as Yucca Mountain (Nevada, United States) have shown that water can migrate rapidly through localized preferential pathways along fractures even under unsaturated conditions (Pruess, 1998; Pruess et al., 1999; Bodvarsson et al., 2003). Laboratory experiments have also illustrated that flow can proceed with minimal imbibition into the surrounding rock matrix, thereby limiting fracture–matrix hydraulic exchange and enabling fast fracture-dominated transport (Liu et al., 1998; Wood et al., 2002; Ji et al., 2004; Tokunaga et al., 2005; Zhou et al., 2022). Furthermore, the geometry and aperture distribution of fracture networks strongly influence recharge efficiency and the rate at which water percolates through the unsaturated zone (Birkholzer et al., 1999; Liu et al., 2002; Cey et al., 2006; Sherman et al., 2020; Zhu et al., 2021).
@@ 3.1 Emergent two-branch retention behavior
$p_{c}(S_{e})=\begin{cases}p_{cc}\left(\dfrac{S_{e}}{S_{c}}\right)^{-1/n_{m}},&S_{e}\leq S_{c}\\[7.5pt] p_{cc}\left(\dfrac{S_{e}-S_{c}}{1-S_{c}}+\varepsilon\right)^{-1/n_{f}}\varepsilon^{1/n_{f}},&S_{e}>S_{c}\end{cases}$ (12)
where $k_{rc}$ is the relative permeability at $S_{c}$ and $p_{cc}$ is the capillary pressure at $S_{c}$. The parameter $\varepsilon$ is introduced to ensure continuity and to regularize the singular behavior at $S_{e}=S_{c}$, defined as:
$\varepsilon=\frac{1}{\left(\dfrac{\alpha_{f}p_{cc}}{\rho g}\right)^{n_{f}}-1}$ (13)
This generalized formulation extends the classical Brooks-Corey relationship by introducing regime-dependent normalization and incorporating the critical saturation (associated with $k_{rc}$ and $p_{cc}$) separating the two regimes. If the effective saturation $S_{e}$ is below the critical saturation $S_{c}$, the system resides in the matrix-dominated regime, in which the constitutive relationships are expressed in terms of the normalized saturation $S_{e}/S_{c}$. For $S_{e}>S_{c}$, corresponding to the fracture-dominated regime, where the constitutive relationships are expressed in terms of the normalized critical saturation $(S_{e}-S_{c})/(1-S_{c})$. The constitutive parameters $n$ and $l$ are assigned with matrix properties for $S_{e}$ below $S_{c}$, and assigned with fracture properties for $S_{e}$ above $S_{c}$. In the limits $S_{c}\to 0$ and $S_{c}\to 1$, the generalized formulation reduces to the classical Brooks–Corey model for fracture-only and matrix-only flow, respectively. This two-branch analytical formulation closely matches the numerical simulation results and well captures the regime transition at $S_{c}$ (Figure 3).
@@ 3.2 Transition between matrix- and fracture-dominated regimes
$h_{c}=\left[\frac{k_{f}\int A_{f}\,d}{k_{m}\int A_{m}\,d}\frac{\alpha_{m}^{2+n_{m}(l_{m}+2)}}{\alpha_{f}^{2+n_{f}(l_{f}+2)}}\right]^{\frac{1}{n_{f}(l_{f}+2)-n_{m}(l_{m}+2)}}$ (16)
@@
Here, we use the approximation $(1-b\gamma)\approx 1$ given that $b\gamma\ll 1$. Combining Equations 9 and 17, and invoking the cubic law for fracture permeability, the critical pressure head $h_{c}$ can be expressed explicitly as a function of fracture aperture:
@@ 4.2 Control of fracture network connectivity
The two-branch retention behavior persists across fracture networks with varying connectivity. Re-examination of Equation 17 shows that the critical transition is controlled primarily by the fracture intensity. The percolation parameter quantifying fracture network connectivity does not explicitly appear in the formulation, but it influences the validity of the analytical prediction. As shown in Figure 4, the analytical solution systematically underestimates the dimensionless critical pressure head for disconnected networks. This discrepancy arises from the assumption in Equation 17 that fractures form fully connected pathways for flow. While this assumption may hold for well-connected networks, where multiple flow paths distribute water flux, it breaks down near or below the percolation threshold (Hyman et al., 2019; Davy et al., 2024). In disconnected networks, fractures alone cannot transmit flow through the system, and a substantial fraction of the flux must pass through matrix, thereby involving fracture–matrix exchange (Yoon et al., 2023), consistent with the limited increase in relative permeability between the critical saturation and full saturation (Text S2 and Figure S8 in Supporting Information). The fracture intensity used in Equation 17 therefore overestimates the hydraulically active fracture population in disconnected networks, leading to an underestimation of the dimensionless critical pressure head.
@@ Acknowledgments
The datasets generated in this study, together with the scripts used for data analysis and figure production, are available in Andiva et al. (2026).